OpenStudy (anonymous):

Is the following equation true? 4 log 3 x + one-half log 3 y – 2 log 3 z = log 3 x to the power of 4 times square root of y all over z-squared. Use complete sentences to explain the properties used in making your decision.

OpenStudy (kirbykirby):

\[4\log_3 x+\frac{1}{2}\log_3y-2\log_3z=\frac{(\log_3x)^4\sqrt{y}}{z^2}\] ??

OpenStudy (kirbykirby):

maybe you can use the Draw tool below if you don't know LaTeX... the equation written like that is just not very clear.

OpenStudy (anonymous):

yes that's what it looks like but the last part is